OpenStudy (xapproachesinfinity):

hello guys, here somewhat challenging substitution integral. \[\int \frac{\sqrt[4]{x}}{\sqrt{x}+\sqrt[3]{x}}dx\]

OpenStudy (solomonzelman):

is it an addition of x^1/3 's on the bottom?

OpenStudy (xapproachesinfinity):

oh sorry the bottom is sqrt(x)+root3(x)

OpenStudy (xapproachesinfinity):

i corrected it

OpenStudy (xapproachesinfinity):

this was in the u sub section

OpenStudy (xapproachesinfinity):

i tried u=x^1/12 and then i had to do another sub along the lines

OpenStudy (xapproachesinfinity):

somehow i arrived at this expression \[\frac{4}{3}x^{1/2}-4x^{1/6}+\arctan(x^{1/2})+c\]

OpenStudy (xapproachesinfinity):

but it is not right i must of did something wrong along the process

OpenStudy (xapproachesinfinity):

http://www.wolframalpha.com/input/?i=int%28%28x%5E1%2F4%29%2F%28x%5E1%2F2%2Bx%5E1%2F3%29%29dx this is wolfram's answer

OpenStudy (kainui):

My process was this, \[\Large x= u^{12} \\ \Large dx = 12 u^{11} du\]

OpenStudy (xapproachesinfinity):

eh that's pretty close to what i did @Kainui i must of just missed the detail perhaps

OpenStudy (kainui):

Let's see, I didn't finish but it looks like this will take us where we need to go.

OpenStudy (xapproachesinfinity):

yes seems so i did u=x^1/12 which is the same as x=u^12

OpenStudy (kainui):

\[\Large \int\limits \frac{u^{12/4}}{u^{12/2}+u^{12/3}} 12 u^{11}du = \\ \Large 12 \int\limits \frac{u^{10}}{1+u^2}du\]

OpenStudy (xapproachesinfinity):

hmm i guess that's where i missed it i got a different expression, darn it lol

OpenStudy (anonymous):

my proccess was this \(\Huge\sf{^P_B{^e_u{^a_t{^n_t{^u_e{^t_r}}}}}\hspace{-65pt}\LaTeX}\)

OpenStudy (anonymous):

sorry... urge was to strong... *runs for the hills*

OpenStudy (kainui):

Yeah, maybe something with exponent rules? \[\Large u=\tan \theta \\ \Large du = \sec^2 \theta d \theta\] \[\Large 12 \int\limits \tan^{10} \theta d \theta\]

OpenStudy (xapproachesinfinity):

yes i checked it i made a mistake with exponents

OpenStudy (xapproachesinfinity):

for me i would do long division for that one?

OpenStudy (xapproachesinfinity):

instead of trig sub?

OpenStudy (xapproachesinfinity):

hmm does not seem a good thing to do long division haha

OpenStudy (xapproachesinfinity):

seems that i got \[12\int( u^8-u^6+u^4-u^2-\frac{1}{1+u^2})du\]

OpenStudy (kainui):

Yeah I mean whatever you gotta do to get the answer haha.

OpenStudy (xapproachesinfinity):

thanks guys i will check my steps and see :)

OpenStudy (anonymous):

what about my solution? you said whatever i got to do to get the answer... xD

OpenStudy (xapproachesinfinity):

by the way how would you deal with tan^10x seems to me it has a lot of work itself

OpenStudy (kainui):

Well you're right about that, I just know there are reduction formulas you could use. I'm not really saying that it's pretty haha.

OpenStudy (xapproachesinfinity):

haha yeah it is not pretty at all seems what i got it correct above i just completed the rest